REACTION OF NFCl_{2} WITH ARGON METASTABLES

\underline{YEVGENIYA I. GERHANOVICH}, JULANNA G. GILBERT, Department of Chemistry, University of Denver, Denver, CO 80208.

The reaction of NFCl_{2} with metastable argon atoms was investigated. When NFCl_{2} was admitted to a stream of argon metastables (^{3}P_{2}), visible and UV emission was observed. Two sources of the visible emission were identified. A sharp peak at 528 nm was assigned to the Δv=0 sequence band of the b^{1}Σ^{+} →X^{3}Σ^{−} transition in NF, and a peak at 664 nm was assigned to the Δv=0 sequence band of the b^{1}Σ^{+} →X^{3}Σ^{−} transition in NCl. The Δv= -1, +1, +2, and +3 sequence bands of the NCl b→X transition were also observed with up to v′=12 levels populated. The vibrational excitation in the NF and NCl fragments has implications for the geometry of the excited state of NFCl_{2} formed in the Ar(^{3}P_{2})+NFCl_{2} reaction . The UV emission spectrum consisted of two bands with maxima at 256 nm and 198 nm. The intensity distribution of the UV emission was identical to the emission that was observed when Cl_{2} was added to the stream of argon metastables and was assigned to electronically excited Cl_{2} formed via Cl_{2} +Ar(^{3}P_{2}). The source of the Cl_{2} is either Cl_{2} contamination in the NFCl_{2} sample, or Cl_{2} formed during the reaction of NFCl_{2} with Ar(^{3}P_{2}). To examine the energetics of the NFCl_{2}+Ar(^{3}P_{2}) reaction, G1 and G2 methods were used to calculate the ΔH_{form} of NFCl_{2}.